Weekly LIPAc highlights (8/March-12/March)

Global Alignment survey

More details on the realignment of the second solenoid magnet (SOL2) of the LEBT (Low Energy Beam Transport) to the entrance of the RFQ completed last week: the achieved tolerance is within 0.1 mm. It is planned to analyse the effects of this change and the measured RFQ-MEBT misalignment by beam dynamics simulation. This will provide key information for the realignment of RFQ-MEBT during the summer maintenance.


This week the injector vacuum was resumed and insulation test for chopper were completed successfully up to 10 kV. The PPS test for starting plasma and HV conditioning were carried out beginning of the week. HV conditioning of the beam source was completed up to 100kV.

Proton extraction started with the initial main scope to confirm the beam characteristics with the re-installed 6mm plasma electrode.

Screenshot of the injector oscilloscope during this week operation with 6mm-PE
  • Duty cycle: 10% (10ms/100ms)
  • Red: Total extracted current (I_ext): 40mA
  • Yellow: Faraday cup signal (I_FC)
  • Blue: Current of repeller PS (I_repeller)
  • Green: Voltage of repeller PS (V_repeller)

Radio Frequency (RF) System and RFQ cavity                                                             

The start-up of the RF system with the calibration of the LLRF (Low Level RF) and the RF chains continued. The check of the RF noise issue on the RF system components continued all the week by increasing the pulse length. The newly installed PYS-LLRF interface to improve the time response of the fast interlock is also improved to cope with spurious interlock generated by the noise issues. During RF injection trials into the RFQ up to 8ms at 1Hz and 128kV were achieved. The RFQ cavity behaviour was good (no instabilities, no interlocks triggered) and up to 4-5% duty cycle no effect was observed on the couplers temperatures. The planned activities to integrate the new RF coupler temperature monitor into the LCS started (remote sessions with INFN are foreseen).

Beam Dump

As part of the BD skid commissioning, the main pump of the BD skid is now operative in manual mode. Circulation in the loop up to the vault and the decay coil was started. It has to be highlighted the remarkable and strenuous effort provided by CIEMAT experts to support remotely the BD skid commission and for solving the LCS and PLC issues.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.